OpenStudy (anonymous):

Claudia works as a manufacturer of decorative jewelry boxes. A customer would like her to create some boxes with the following conditions: The sum of the length and width must be 30 centimeters. The height must be 3 centimeters less than the length. The box should have the greatest possible volume. Let's do some math to see if we can decide what size boxes Claudia needs to make, and show what we know about polynomials in the process. Remember: One formula we can use to find the volume of a box is If the length of the box is x, write an expression for the width in terms of x.

OpenStudy (anonymous):

I have 8 more questions that go along with this one.... Please helpppp

OpenStudy (superdavesuper):

"length of the box is x" "sum of the length and width must be 30 centimeters" so what is the width in terms of x? Take a guess plz?

OpenStudy (anonymous):

(x)(l)=30?????

OpenStudy (superdavesuper):

Sum means adding...

OpenStudy (anonymous):

Oh wow... okay so x+l=30

OpenStudy (superdavesuper):

Yes, so the width = ?

OpenStudy (anonymous):

x..... 30/l???

OpenStudy (superdavesuper):

remember u wrote x + width = 30 right? so....

OpenStudy (anonymous):

30-x=width

OpenStudy (superdavesuper):

That is the right answer :)

OpenStudy (anonymous):

Ooh okay so now... If the length of the box is x, write an expression for the height in terms of x.

OpenStudy (superdavesuper):

The height must be 3 centimeters less than the length. So...

OpenStudy (anonymous):

length-3= height ?

OpenStudy (anonymous):

So x-3=height

OpenStudy (superdavesuper):

That is correct :)

OpenStudy (anonymous):

Ah thank you. Ok only 7 more... Write a function f(x) that gives the volume of a box of length x in polynomial form.

OpenStudy (superdavesuper):

Sorry I got 5 more min only. Is the box a cube?

OpenStudy (anonymous):

Oh no... okay well it doesn't say.

OpenStudy (superdavesuper):

Hmmm just length by itself is no good to calculate volume. sorry!

OpenStudy (anonymous):

it says V=lwh

OpenStudy (superdavesuper):

yes but u only get x...oh wait - is this part of the same question in the first post?

OpenStudy (anonymous):

yes! (:

OpenStudy (superdavesuper):

OK then use the width n height eqns from above and put them in V=lwh :)

OpenStudy (anonymous):

(30-x)(x-3) ?

OpenStudy (superdavesuper):

one more term....

OpenStudy (anonymous):

wait is it: (30-width)(30-x)(x-3)

OpenStudy (superdavesuper):

Where u get (x-3)?? x = length right?

OpenStudy (superdavesuper):

No no no ignore the last one....

OpenStudy (superdavesuper):

(30-x)(x-3) and u missed the length....

OpenStudy (anonymous):

x is length... so (30-width)

OpenStudy (anonymous):

So (30-width) (30-x) (x-3) (l)(w)(h)

OpenStudy (superdavesuper):

express length in term of x pl

OpenStudy (superdavesuper):

u said "x is length... "

OpenStudy (anonymous):

Yeah... I think I am. 30-width=length. 30-x=width x-3=height. The x is length

OpenStudy (superdavesuper):

"x is length... " "30-x=width" "x-3=height"

OpenStudy (anonymous):

Im confused.... 30-x=width is 30-length=width, 30-width=x which is 30-width=length and x-3=height is length-3=height.

OpenStudy (superdavesuper):

Put length, width n height ALL in terms of x pl

OpenStudy (anonymous):

(30-x=w) (30-w=x) (x-3=h)

OpenStudy (superdavesuper):

U want the V=lwh equation with ONLY x, NO other unknown ok?

OpenStudy (superdavesuper):

no l, no w, no h, no =, nothing but x and numbers :)

OpenStudy (anonymous):

Ok so how do I know what numbers to put in replace of w and h?

OpenStudy (superdavesuper):

"x is length... " "30-x=width" "x-3=height"

OpenStudy (anonymous):

So I choose the numbers randomly??

OpenStudy (superdavesuper):

I have to go; here is it V=x(30-x)(x-3)

OpenStudy (anonymous):

Oh okay. Thank you very much for your help!!!!!!! (: